OpenStudy (anonymous):

Using the chain rule, find the derivative of: (4x+3)^4 * (x+1)^-3

OpenStudy (anonymous):

fine here...

OpenStudy (anonymous):

use quotient rule... then apply u substitution as you go

OpenStudy (anonymous):

\[(f * g' - f' * g) / g^2\]

OpenStudy (anonymous):

\[(4x+3)^4 * 3(x+1)^3 - 16(4x + 3)^4 * (x+1)^3 / (x+1)^6\]

OpenStudy (anonymous):

u get 3(x+1)^3 from..... u = x+1, du = 1

OpenStudy (anonymous):

(x+1)^3 then becomes (u)^3... which the derivative is then 3 * (u)^2

OpenStudy (anonymous):

er messed up 2 steps up.... it's supposed to be 3(x+1)^2 and 16(4x+3)^3

OpenStudy (anonymous):

then substitute the u = x+1 into 3 * (u)^2 and it becomes 3(x+1)^2

OpenStudy (anonymous):

do the exact same steps for f' => u = 4x+3, du = 4

OpenStudy (anonymous):

then just cancel out the (x+1)^6 with the (x+1)'s in the numerator. you don't have to split the fraction into two but you can

OpenStudy (anonymous):

answer then becomes

OpenStudy (anonymous):

\[(3(4x+3)^4 * (x+1)^2- 16(4x+3)^3 * (x+1)^3 )/ (x+1)^6\]

OpenStudy (anonymous):

\[(3(4x+3)4∗(x+1)2) / (x+1)^6 −(16(4x+3)3∗(x+1)3)/(x+1)^6\]

OpenStudy (anonymous):

All i have so far understood is u = (4x+3)^4 u' = 4(4x+3)^3 v = (x+1)^3 v' = 3(x+1)^2 vu' - uv' / v²

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

u = (4x+3) not u = (4x+3)^4

OpenStudy (anonymous):

you do this to simplify taking a derivative, so if you were to set u = 4x+3 then the original (4x+3)^4 becomes (u)^4

OpenStudy (anonymous):

for the sake of the quotient rule first

OpenStudy (anonymous):

we havent even gotten to the chain rule part yet

OpenStudy (anonymous):

you did the chain rule with the 4(4x+3)^3

OpenStudy (anonymous):

correct me if i'm wrong.... lets say u = (4x+3)^4 u' = 4(4x+3)^3 and lets use....w as the derivative of the inside so w = 4 then for the bottom of the quotient v = (x+1)^3 v' = 3(x+1)^2 and x for the derivative of the inside so x = 1

OpenStudy (anonymous):

now vu'w - uv'x / v²

OpenStudy (anonymous):

right however you don't do u substitution in that way

OpenStudy (anonymous):

here the idea for this is you can use u and du rather than w or x or whatever

OpenStudy (anonymous):

so when you say u = (4x+3)^4... that is an incorrect way of doing it

OpenStudy (anonymous):

i understand that but the quotient rule teaches us that vu' - uv' / v²

OpenStudy (anonymous):

what you do is you set u = 4x+3 then du = 4 (not w.. du means the derivative of u)

OpenStudy (anonymous):

i understand that...there's only so many letters in the alphabet

OpenStudy (anonymous):

so then you have (u)^4 * du

OpenStudy (anonymous):

er sorry the derivative with respect to u and using the chain rule (u)^3 * du

OpenStudy (anonymous):

man this is making no sense

OpenStudy (anonymous):

yeah but you have to do u and du... right now it seems trivial but the u and du system will be used in integrals and other stuff which requires the u and du system

OpenStudy (anonymous):

START OVER (4x+3)^4 / (x+1)^3

OpenStudy (anonymous):

yeah u substitution is kind of tricky at first

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

step 1: i wrote the problem step 2: ?

OpenStudy (anonymous):

ok so basically what you do is you use the quotient rule just like you normally would

OpenStudy (anonymous):

so uv' - u'v / v^2

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

er u'v - uv' / v^2

OpenStudy (anonymous):

(x+1)^3 (4)(4x+3)^3 - 3(x+1)²(4x+3)^4 / (x+1)^6

OpenStudy (anonymous):

remember to multiply by du as well.. so on the (4x+3)^4, you brought down the 4 which is 4(4x+3)^3 but you also have to multiply by du which = 4. it becomes 16(4x+3)^3

OpenStudy (anonymous):

(4)(x+1)^3 (4)(4x+3)^3 - (1)3(x+1)²(4x+3)^4 / (x+1)^6 ??? 4 for the derivative of 4x + 3 1 for the derivative of x+1

OpenStudy (anonymous):

right 4 for the derivative but you also brought down the 4 from the exponent

OpenStudy (anonymous):

(x+1)^3 (16)(4x+3)^3) - 3(x+1)²(4x+3)^4 / (x+1)^6

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

u got it

OpenStudy (anonymous):

i showed bringing down the 4 from the exponent but hadnt touched the du of the insides yet

OpenStudy (anonymous):

now how do you simplify?

OpenStudy (anonymous):

k then simplifying is just algebra. so what you can do is split it up into two fractions or you can pull the (x+1)'s out using partial fraction decomposition (fuk this method... way too much work.. use the first)

OpenStudy (anonymous):

\[(16(4x+3)^3∗(x+1)^3)/(x+1)^6-(3(4x+3)^4∗(x+1)^2)/(x+1)^6\]

OpenStudy (anonymous):

sec let me check this

OpenStudy (anonymous):

k yeh its right

OpenStudy (anonymous):

so you basically just split up the equation

OpenStudy (anonymous):

(16(4x+3)^3∗(x+1)^3)/(x+1)^6 can be (16(4x+3)^3∗(x+1)^3)/ (x+1)^3 (x+1)^3 that splits the denominator into 2 terms of (x+1)^3........... one of those will cancel the (x+1)^3 on top

OpenStudy (anonymous):

right so after canceling out you get (x+1)^3 on the bottom

OpenStudy (anonymous):

then do the same for the other part

OpenStudy (anonymous):

answer in the back of the book: (4x+3)^3(4x+7) / (x+1)^4

OpenStudy (anonymous):

i get 3(4x+3)^4 / (x+1)^4 for the other part

OpenStudy (anonymous):

no idea where the (4x+7) term comes from...it is the same answer though

OpenStudy (anonymous):

the answer in the back of the book is the same answer. The book used partial fraction decomposition though... which is a pain in the retrice

OpenStudy (anonymous):

i think partial fraction was algebra.. dunno you might as well learn it because you will have to use it later in calculus once or twice, but it's really not that useful in my eyes

OpenStudy (anonymous):

here check this stuff out on youtube. patrickjmt = calculus god.

OpenStudy (anonymous):

chain rule: http://www.youtube.com/watch?v=6kScLENCXLg

OpenStudy (anonymous):

hrm can't find any u substitution from him except for integrals, which you will learn closer to your final. http://www.youtube.com/watch?v=qclrs-1rpKI

OpenStudy (anonymous):

ok so i'm starting this other problem now.....this one uses the product rule and makes more sense... y = sin(x²)cos(2x) so i'm gonna do uv' d(2x) + vu' d(x²)

OpenStudy (anonymous):

they may have not taught you u-sub yet, but i highly recommended getting a good understanding of it early. (it's really not that difficult anyway. u just need a quick vid tutorial on it)

OpenStudy (anonymous):

so my answer will be sin(x²)(-sin(2x))(2) + (2x)(cos(2x))(cos(x²)

OpenStudy (anonymous):

yeh, just do the same thing. plug and chug. let u = x^2, du = 2x. take derivative of sinu, which = cosu. multiply by du. 2xcos(x^2)

OpenStudy (anonymous):

also written as -2sin(x²)sin(2x) + 2xcos(2x)cos(x²)

OpenStudy (anonymous):

your answer is right. remember your trig identities though

OpenStudy (anonymous):

but it's right

OpenStudy (anonymous):

wtf ya'll are doing dot product already? we just finished dot product and i'm in cal 3

OpenStudy (anonymous):

dot product?

OpenStudy (anonymous):

nm

OpenStudy (anonymous):

k got some studying to finish up and you pretty much have it. watch the patrickJMT calculus videos. he's a beast at teaching